2. martin corporation's common stock sells for $55 per share. the current dividend is $2.75 per share; dividends are expected to grow at 5.00 percent per year indefinitely. the floatation cost of new common stock is 20.00 percent. what is the firm's cost retained earnings?
Answer #1

COST OF RETAINED EARNINGS : Do not include flotation costs

Cost of retained earnings = re = D1/P0 + g

D0 = 2.25, g = 5%, P0 = 55

D1 = D0*(1+g) = 2.75 *(1+0.05) = 2.8875

re = 2.8875/55 + 0.05 = 0.1025 = 10.25%

Answer : 10.25% (Thumbs up please)